OpenStudy (anonymous):

Use th method of substitution to solve the system of linear equation.(numbr three key not working..) 2x+5y= two/ thirds y=three x-1 The solution is?

OpenStudy (anonymous):

Substitution means substituting one variable with another so that there is only one variable (in, this case, x) in the equation and it can be solved for. The first step is putting y in terms of x. So: 2x + 5(3x-1)=2/3 Can you solve it from here?

OpenStudy (anonymous):

ok i will try

OpenStudy (anonymous):

2x+15x-5=two/third 17x-5=2/third

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

sorry, I'm here. once sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

17x-5=2/3 Okay, so now you still have to isolate x. Do you know how to do that? To get x alone on one side of the equation, with everything else on the other side?

OpenStudy (anonymous):

do i add 5 to both sids and get 17/7

OpenStudy (anonymous):

5+2=7

OpenStudy (anonymous):

y=7x-17

OpenStudy (anonymous):

Wait, go back!

OpenStudy (anonymous):

?

OpenStudy (anonymous):

where did i go wrong?

OpenStudy (anonymous):

you lost me at y= 7x- 17, but 5 + 2/3 doesn't equal 17/7

OpenStudy (anonymous):

ok this is the part where i get confused

OpenStudy (anonymous):

So, from there: 17x-5=2/3 17x= 5 + 2/3 (or 17/3)

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

ok so i have to work the other stp to solve the 2nd solution

OpenStudy (anonymous):

for y

OpenStudy (anonymous):

y=17x-three

OpenStudy (anonymous):

Wait, you still haven't isolated x though!

OpenStudy (anonymous):

ok i skipt ahead

OpenStudy (anonymous):

ok lets solve

OpenStudy (anonymous):

you have one more step in order to get x by itself. 17x=17/3 Now divide both sides by 17 (or, to make it easier when multiplying fractions, you can multiply by 1/17 x = 17/3 * 1/17

OpenStudy (anonymous):

so x=1/3

OpenStudy (anonymous):

NOW you can plug x into y=3x -1

OpenStudy (anonymous):

ok so y=three x-1 y= three*1/thid*1 y=0

OpenStudy (anonymous):

yes!! awesome! Now to check it: 2(1/3)+5(0)= 2/3 2/3 + 0 = 2/3 nice work!

OpenStudy (anonymous):

so the solution is (2/third,0)

OpenStudy (anonymous):

im sorry i am rally trying please bare with me

OpenStudy (anonymous):

really

OpenStudy (anonymous):

you're doing well! And yes, the solution is x=2/3, y=0 or (2/3,0)

OpenStudy (anonymous):

ok i am trying so hard i really do appreciate your help

OpenStudy (anonymous):

I used to be TERRIBLE at math, the only reason I got better is because my brother worked with me for an hour every single freaking day. stick with it! do you have any other questions?

OpenStudy (anonymous):

no i am adding you